How would I multiply (4+a)^2? This is what I have 16+a. Am I on the correct path?
Found 2 solutions by rfer, stanbon:
Answer by rfer(16322)

(4+a)(4+a)
16+4a+4a+a^2=
16+8a+a^2

Answer by stanbon(75887)

How would I multiply (4+a)^2?
This is what I have 16+a.
Am I on the correct path?
-----
(4+a)(4+a) = 4(4+a)+a(4+a) = 16 + 4a + 4a + a^2 = a^2 + 8a + 16
--------------------
Cheers,
Stan H.
------------